OpenStudy (twistrose):

2. The average distance from the sun to the planet Mercury is about 58,000,000 km. The diameter of a human hair is about 0.0025 cm. (a) What is the distance from the sun to Mercury written in scientific notation? (b) What is the diameter of a human hair written in scientific notation? (c) Fill in the blank: When writing a number in scientific notation, the coefficient has to be a number that is greater than or equal to 1 but still be less than ______. (d) True or False. The exponent indicates how many places to move the decimal and in which direction.

OpenStudy (jango_in_dtown):

so @Twistrose you tried? any answers?

OpenStudy (mhchen):

any ideas?

OpenStudy (twistrose):

yes i tried but it mostly got me at scientific notation, which i cant do because i have a hard time figuring it out. maybe if i had that you think i could get he others?

OpenStudy (jango_in_dtown):

ok... you must have a*10^n, where a is any real number more than 1 but less than 10

OpenStudy (jango_in_dtown):

and n is an integer

OpenStudy (jango_in_dtown):

58,000,000=5.8 *10^ ??

OpenStudy (jango_in_dtown):

0.0025=2.5*10^

OpenStudy (jango_in_dtown):

(c) answe is 10

OpenStudy (jango_in_dtown):

d is true

OpenStudy (jango_in_dtown):

a and b you try. I gave hints

OpenStudy (twistrose):

well for the problem you gave me i got 336,400,000 is that right

OpenStudy (twistrose):

so thst is 3.3*10^5 right or did i do it wrong

OpenStudy (rcp031):

scientific notation? means is a special way of writing numbers:

OpenStudy (rcp031):

How to Do it To figure out the power of 10, think "how many places do I move the decimal point?" When the number is 10 or greater, the decimal point has to move to the left, and the power of 10 is positive. When the number is smaller than 1, the decimal point has to move to the right, so the power of 10 is negative. Example: 0.0055 is written 5.5 × 10-3 Because 0.0055 = 5.5 × 0.001 = 5.5 × 10-3 Example: 3.2 is written 3.2 × 100 We didn't have to move the decimal point at all, so the power is 100 But it is now in Scientific Notation